OpenStudy (anonymous):

What two numbers have the product of 234 and the sum of 31

OpenStudy (danjs):

the product of 2, 3, and 4? can you elaborate a little bit more please. and Welcome to open study

OpenStudy (danjs):

I think the question might be typed in a bit wrong, not sure , can you check on that please

OpenStudy (anonymous):

No I just checked and couldn't get the answer

OpenStudy (danjs):

ok, i see what they want.

OpenStudy (danjs):

define two variables for your two unknown numbers, let them be X and Y The sum of the two numbers x and y are 31 X + Y = 31 The Product of x and y is 234 X * Y = 234

OpenStudy (danjs):

You now have 2 equations, with two unknown variables, it is possible to solve these for X and Y

OpenStudy (danjs):

X + Y = 31 X * Y = 234

OpenStudy (danjs):

Do you know where to begin here?

OpenStudy (anonymous):

Yes that's the equation but they need to have the same 2 numbers and that's whats I cant get

OpenStudy (danjs):

right, we need to find a value for x and y, so that both those equations are true

OpenStudy (danjs):

to start, solve X+ Y = 31 for X to get X = 31 - Y

OpenStudy (danjs):

Now you can substitute that X = 31 - Y into the second equation like this X*Y = 234 (31 - Y)*Y = 234 31Y - Y^2 = 234 Y^2 - 31Y + 234 = 0 (Y-18)(Y-13) = 0

OpenStudy (danjs):

i distributed the left side, and then moved it all to one side and factored into the two parenthesis terms, So now you have (Y-18)(Y-13)=0 which means either one of those can be zero Y-18 = 0 Y - 13 = 0 Y = 13 OR Y=18

OpenStudy (anonymous):

Oh my gosh thank you so much!!!

OpenStudy (danjs):

If Y=13 Then from equation 1, X+ Y = 31 X + 13 = 31 X = 31 - 13 = 18

OpenStudy (danjs):

SO the choices are X=13, Y=18 Or X=18 , Y = 13

OpenStudy (danjs):

The two numbers are 13 and 18

OpenStudy (danjs):

U understand how the equations were pulled from the problem X + Y = 31 X*Y = 234

OpenStudy (danjs):

If you would like, hover over my name and click ADD FAN, then i will be in your friends list at the top so you can as me questions later if you want. In the question type @DanJS

OpenStudy (anonymous):

Thank your i did

OpenStudy (anonymous):

You

OpenStudy (zzr0ck3r):

another way 234=2*9*13=18*13